# TritonRL: Training LLMs to Think and Code Triton Without Cheating

### v1

@@ 2.2 Reinforcement Learning with Hierarchical Reward Decomposition
$\begin{aligned} \mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim P(Q),~\{o_{i}=(o_{i}^{\text{plan}},o_{i}^{\text{code}})\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\left[\frac{1}{G}\sum_{i=1}^{G}\alpha\mathcal{F}_{\text{GRPO}}^{\text{plan}}(\theta,i)+\mathcal{F}_{\text{GRPO}}^{\text{code}}(\theta,i)\right]\end{aligned}$
Figure 3: An example LLM output for Triton code generation, showing a reasoning trace (plan) and the generated Triton kernel code conditioned on the plan.
where $\pi_{\theta}$ and $\pi_{\theta_{old}}$ are the policy model and reference model, $q$ denotes a prompt given to the model, defining a task to implement in Triton, and $o_{i}=(o_{i}^{\text{plan}},o_{i}^{\text{code}})$ represents $i$-th response generated by the model for $q$ in the group $G$. Here we denote $o_{i,t}^{\text{plan}}$ and $o_{i,t}^{\text{code}}$ as $t$-th token of output plan and code, and $T_{i}^{\text{plan}}$ and $T_{i}^{\text{code}}$ as the sets of token positions in the plan and code, respectively. The GRPO losses $\mathcal{F}^{\text{plan}}(\theta)$ and $\mathcal{F}^{\text{code}}(\theta)$ are computed over the tokens in generated plans and Triton codes, respectively, and $\alpha\in[0,1]$ is a weighting factor that balances the training speed of planning and coding policy. Here, $\alpha$ is set to a small value (e.g., $0.1$) so that the planning distribution is updated slowly, allowing the coding policy sufficient time to learn correct implementations conditioned on those high-level plans. The detailed formulation of each loss component is as follows.
@@
$\begin{aligned} \mathcal{F}_{\text{GRPO}}^{\text{plan}}&=\frac{1}{|o_{i}^{\text{plan}}|}\sum_{t\in T_{i}^{\text{plan}}}A_{i}^{\text{plan}}\cdot\min\left\{\frac{\pi_{\theta}(o_{i,t}^{\text{plan}}|q,o_{i,<t}^{\text{plan}})}{\pi_{\theta_{old}}(o_{i,t}^{\text{plan}}|q,o_{i,<t}^{\text{plan}})},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}^{\text{plan}}|q,o_{i,<t}^{\text{plan}})}{\pi_{\theta_{old}}(o_{i,t}^{\text{plan}}|q,o_{i,<t}^{\text{plan}})},1-\epsilon,1+\epsilon\right)\right\},\\ \mathcal{F}_{\text{GRPO}}^{\text{code}}&=\frac{1}{|o_{i}^{\text{code}}|}\sum_{t\in T_{i}^{\text{code}}}A_{i}^{\text{code}}\cdot\min\left\{\frac{\pi_{\theta}(o_{i,t}^{\text{code}}|q,o_{i}^{\text{plan}},o_{i,<t}^{\text{code}})}{\pi_{\theta_{old}}(o_{i,t}^{\text{code}}|q,o_{i}^{\text{plan}},o_{i,<t}^{\text{code}})},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}^{\text{code}}|q,o_{i}^{\text{plan}},o_{i,<t}^{\text{code}})}{\pi_{\theta_{old}}(o_{i,t}^{\text{code}}|q,o_{i}^{\text{plan}},o_{i,<t}^{\text{code}})},1-\epsilon,1+\epsilon\right)\right\},\end{aligned}$
where $A_{i}^{\text{plan}}$ and $A_{i}^{\text{code}}$ are the group-wise advantages for plan and code tokens, computed as $A_{i}=r_{i}-\frac{1}{G}\sum_{j=1}^{G}r_{j}$, with rewards for plan and code tokens defined as:
$r_{i}^{\text{plan}}=\texttt{syntax}_{i}\cdot\texttt{func}_{i}\cdot\texttt{speedup}_{i},\quad r_{i}^{\text{code}}=\texttt{syntax}_{i}\cdot\texttt{func}_{i}\cdot\texttt{correct}_{i}.$ (3)
Here, we note that the syntax and functionality checks serve as necessary conditions for correctness and speedup evaluations. If either the syntax or functionality check fails, the generated code is deemed invalid, and both correctness and speedup rewards are set to zero.
@@
To provide separate feedback for the hierarchical actions of planning and coding with their respective contributions, we design distinct reward functions for each action. As described in equation 3, we decouple rewards for high-level plan tokens and low-level code tokens based on their main contributions to the final output, assigning speedup-based rewards to plan tokens and correctness-based rewards to code tokens. This approach ensures that reasoning traces are encouraged to propose optimization strategies that yield efficient kernels, while code generation is guided to produce valid implementations that faithfully realize these plans. In this way, our hierarchical reward decomposition provides a more targeted and effective learning signal compared to uniform reward designs used in prior work Li et al. (2025b); Baronio et al. (2025), which apply a uniform reward to all tokens.
Another key aspect of our reward decomposition is the use of the weighting factor $\alpha$ to balance the training speed of planning and coding policies. To effectively learn both planning and coding skills, it is important to ensure that the model has sufficient time to learn correct Triton implementation given the current planning distribution. If $\alpha$ is set to $1.0$, the planning and coding policies are updated at the same rate, which may lead to instability in learning, as the coding policy may struggle to keep up with rapidly changing plans. In contrast, if $\alpha$ is too small (e.g., $0.0$), the planning policy remains static and only the coding policy is updated, which may limit the model’s ability to explore promising kernel optimization plans. By setting $\alpha$ to an appropriately small value (e.g., $0.1$), we ensure that the planning distribution is updated slowly, allowing the coding policy sufficient time to learn correct implementations conditioned on those high-level plans. This helps avoid overly penalizing the reasoning trace due to code generation instability in early training, thus preserving promising optimization plans that may yield higher speedup once the code generation stabilizes.
In our experiments, we compare our hierarchical reward decomposition with uniform reward designs explored by prior work Li et al. (2025b); Baronio et al. (2025), and evaluate different choices of $\alpha$ to highlight the advantages of the hierarchical reward decomposition.
Difficulty-Aware Data Mixing.
The selection of training data is crucial for effective RL post-training, and many works have shown that selective sampling by leveraging additional information, such as difficulty or interaction between data points, can significantly improve model performance (Yu et al., 2025; Chen et al., 2025; 2024). While KernelBook Paliskara & Saroufim (2025) is a large-scale dataset containing diverse kernel generation tasks, uniformly sampling from the entire dataset may be inefficient and lead to suboptimal model performance. In this work, we augment each task in KernelBook with a difficulty label and explore different data mixtures to construct an optimal training set for RL post-training.
We use an LLM labeler to create difficulty labels for each data point in KernelBook based on the difficulty levels defined in KernelBench (Ouyang et al., 2025), which categorizes tasks into three levels based on the complexity of kernel implementation, and we denote the subset of training data with difficulty level $d$ as $\mathcal{D}_{train}^{(d)}$ for $d\in\{1,2,3\}$. Forcusing on level 1 and 2 tasks, we form various data mixtures by sampling from these subsets with data mixing probabilities $\bm{p}=(p_{1},p_{2})$, where $p_{d}$ denotes the probability of sampling from $\mathcal{D}_{train}^{(d)}$. By varying the mixture probabilities $\bm{p}$, we can construct different training data mixtures. In our experiments, we explore multiple mixtures and evaluate the performance of the post-trained model on KernelBench to identify the optimal training data configuration. Additional details are provided in Appendix F.1.
@@ 3.1 Training and Evaluation Setups
Data Preparation.
For both SFT and RL, we use 11k tasks from KernelBook (Paliskara & Saroufim, 2025). We expand each task with five reasoning traces and corresponding Triton implementations generated by DeepSeek-R1 (Guo et al., 2025), yielding 58k <task query, Triton code with CoT> pairs. Prompts adopt a one-shot format, where the reference PyTorch code is given and the model is asked to produce an optimized Triton alternative (examples in Appendix F.5). To support curriculum in RL training, we further label tasks into three difficulty levels using Qwen3-235B-Instruct (Team, 2025), following the task definitions in Ouyang et al. (2025), yielding 11k <task query, level> pairs. See detailed generation and classification in Appendix H and  F.1.
@@
We begin by fine-tuning the base model Qwen3-8B on KernelBook tasks. After SFT, we move to RL training on the same KernelBook tasks, but without output labels—rewards are computed directly from code execution-so the RL dataset consists only of task instructions. An example of such instructions is shown in Appendix F.5. We implement training under the VeRL framework (Sheng et al., 2025). Unlike SFT, we use the subset of KernelBook tasks labeled as Level 1 and Level 2 for RL training. We explore different data mixtures from Level 1 and Level 2 tasks for RL training, where the default mixture is $\bm{p}=[1.0,0.0]$ unless specified otherwise. Hyperparameters for both SFT and RL are provided in Appendix H. By default, we use the reward function $r$ from equation 3, setting $\alpha^{*}=0.1$ unless specified otherwise.
@@
We evaluate TritonRL on KernelBench (Li et al., 2025a) [^2] We use the Triton backend version of KernelBench from https://github.com/ScalingIntelligence/KernelBench/pull/35.. KernelBench offers an evaluation framework covering 250 tasks, divided into Level 1 (100 single-kernel tasks, such as convolution), Level 2 (100 simple fusion tasks, such as conv+bias+ReLU), and Level 3 (50 full architecture tasks, such as MobileNet), to assess LLM proficiency in generating efficient CUDA kernels. We conduct experiments mainly on the Level 1 and Level 2 tasks from KernelBench. The prompts used for these benchmarks are provided in Appendix G.1.
@@
We evaluate the performance of LLMs for generating Triton code in terms of (1) Validity (syntax and functionality); (2) Correctness (compilation and correct output); (3) Speedup (relative execution time improvement). We report fast1 and fast2 to indicate the model’s ability to generate Triton code that is at least as fast as or twice as fast as the reference PyTorch implementation, respectively. The formal definition of metrics is provided in Appendix D. We measure the pass@$k$ metrics for each aspect, which indicates the ratio of generating at least one successful solution among $k$ sampled attempts. We use $k=10$ as a default unless specified. We test both Triton codes and reference PyTorch codes on an NVIDIA L40S.
@@
We compare TritonRL with several baselines, including KernelLLM (Fisches et al., 2025) and AutoTriton (Li et al., 2025b), which are fine-tuned LLMs specifically for Triton programming. We also include our base model Qwen3-8B (Team, 2025) without any fine-tuning, fine-tuned Qwen3-8B only after SFT, and larger Qwen3 models (e.g., Qwen3-14B and Qwen3-32B). Additionally, we evaluate Claude-3.7 (Anthropic, 2025) with unknown model size. For large model classes beyond 100B (e.g., GPT-OSS 120B (OpenAI, 2025), DeepSeek-R1-0528 (NVIDIA Developer Blog, 2025)), we report the numbers to the Appendix I.1 as a reference.
Model #Params Level1 (robust verifier) Level1 (w/o robust verifier)
valid compiled / correct $\textbf{fast}_{1}$ / $\textbf{fast}_{2}$ mean speedup valid compiled / correct
Qwen3 (base) 88B $73.0$ $40.0$ / $14.0$ $0.0$ / $0.0$ $0.03$ $54.0$ $52.0$ / $15.0$
Qwen3 1414B $82.0$ $65.0$ / $17.0$ $0.0$ / $0.0$ $0.04$ $66.0$ $71.0$ / $16.0$
Qwen3 3232B $75.0$ $61.0$ / $16.0$ $2.0$ / $0.0$ $0.06$ $52.0$ $50.0$ / $15.0$
KernelLLM 88B $42.0$ $40.0$ / $20.0$ $0.0$ / $0.0$ $0.05$ $100.0$ $98.0$ / $29.0$
AutoTriton 88B $97.0$ $78.0$ / $50.0$ $2.0$ / $1.0$ $0.25$ $100.0$ $95.0$ / $70.0$
TritonRL (ours) 88B $99.0$ $82.0$ / $\mathbf{56.0}$ $\mathbf{5.0}$ / $1.0$ $\mathbf{0.33}$ $99.0$ $83.0$ / $58.0$
   w/o RL (SFT only) 88B $97.0$ $\mathbf{88.0}$ / $44.0$ $4.0$ / $\mathbf{2.0}$ $\mathbf{0.33}$ $98.0$ $93.0$ / $47.0$
Claude-3.7 - $99.0$ $\mathbf{99.0}$ / $53.0$ $3.0$ / $1.0$ $0.32$ $100.0$ $100.0$ / $64.0$
Table 1: Main results on KernelBench Level 1. All metrics are reported as pass@10 (%). Our model achieves the best results among models with fewer than 32B parameters. The left block reports evaluation with the robust verifier (syntax + functionality). The right block (w/o robust verifier) lacks functionality checks, leading to misleading correctness estimates.
Figure 4: Pass@$k$ correctness and mean speedup for $k=1,5,10$ on KernelBench Level 1 tasks. The figures show how performance (correctness on the left, mean speedup on the right) of TritonRL and baseline models scales as the number of attempts increases.

### v2

@@ 2.3 Policy Optimization with Hierarchical Reward Decomposition
$\begin{aligned} \mathcal{J}(\theta)&=\mathbb{E}_{q_{i}\sim\mathcal{D}_{RL},~\{o_{i,j}\}_{j=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q_{i})}\left[\alpha\mathcal{J}_{i}^{\text{\tiny plan}}(\theta)+\mathcal{J}_{i}^{\text{\tiny code}}(\theta)\right]\end{aligned}$
where $\alpha\in[0,1]$ is a weighting factor that balances the update rates between the model’s planning and coding actions. For each action class $c\in\{\text{plan},\text{code}\}$, the class-specific objective $\mathcal{J}_{i}^{c}(\theta)$ is defined as:
@@
$\begin{aligned} \mathcal{J}_{i}^{c}(\theta)=\frac{1}{G}\sum_{j=1}^{G}\frac{1}{|o_{i,j}|}\sum_{t\in T_{i,j}^{c}}\min\Big[&A_{i,j}^{c}s_{i,j,t}^{c}(\theta),\cr&A_{i,j}^{c}\text{clip}\left(s_{i,j,t}^{c}(\theta),1-\epsilon,1+\epsilon\right)\Big],\end{aligned}$
where $T_{i,j}^{c}$ represents the set of token indices for class $c$ in output $o_{i,j}$ and the token-wise probability ratios are given by
$\begin{aligned} s_{i,j,t}^{\text{\tiny plan}}(\theta)&=\frac{\pi_{\theta}(o_{i,j,t}^{\text{\tiny plan}}|q_{i},o_{i,j,<t}^{\text{\tiny plan}})}{\pi_{\theta_{old}}(o_{i,j,t}^{\text{\tiny plan}}|q_{i},o_{i,j,<t}^{\text{\tiny plan}})},\\ s_{i,j,t}^{\text{\tiny code}}(\theta)&=\frac{\pi_{\theta}(o_{i,j,t}^{\text{\tiny code}}|q_{i},o_{i,j}^{\text{\tiny plan}},o_{i,j,<t}^{\text{\tiny code}})}{\pi_{\theta_{old}}(o_{i,j,t}^{\text{\tiny code}}|q_{i},o_{i,j}^{\text{\tiny plan}},o_{i,j,<t}^{\text{\tiny code}})}.\end{aligned}$
The group-wise advantages $A_{i,j}^{c}$ are computed as $A_{i,j}^{c}=r_{i,j}^{c}-\frac{1}{G}\sum_{j=1}^{G}r_{i,j}^{c}$, where we assign different reinforcement signal for each token class as follows:
$r_{i,j}^{\text{\tiny plan}}=R^{\text{speedup}}(g_{i},o_{i,j}),\quad r_{i,j}^{\text{\tiny code}}=R^{\text{correct}}(g_{i},o_{i,j}),$ (2)
where the reward functions are defined in (1).
@@
Hierarchical vs Uniform Rewards. Unlike uniform reward designs (Li et al., 2025b; Baronio et al., 2025) that apply the same signal to all tokens, our HRD provides targeted feedback. In (2), we assign speedup-based rewards to plan tokens to encourage the discovery of efficient optimization strategies, while assigning correctness-based rewards to code tokens. This separation ensures that reasoning traces are encouraged to propose optimization strategies that yield efficient kernels, while code generation is guided to produce valid implementations that faithfully realize these plans. Assigning correctness-based rewards to code tokens is deliberate because the implementation is conditioned on the plan, penalizing code for low speedup would unfairly punish a correct implementation of an inherently suboptimal plan. Correctness rewards for code tokens decouple this confound, allowing the code policy to be reinforced purely for faithful realization of a given plan. We validate this design choice in Appendix F.4, where we compare assigning correctness-based versus speedup-based rewards to code tokens, supporting our rationale.
Balancing Planning and Coding Updates. The weighting factor $\alpha$ manages the co-evolution of the planning and coding policies. To effectively master both, the model requires a stable ”practice environment.” If $\alpha=1.0$, the planning distribution shifts too rapidly for the coding policy to stabilize. By setting $\alpha\approx 0.1$, we ensure the planning distribution evolves gradually, giving the coding policy sufficient ”time” to learn correct implementations for a given set of plans. This avoids the premature rejection of promising optimization strategies due to transient implementation errors in early training phases.
Table 1: Main results on KernelBench Level 1 and Level 2. All metrics are reported as pass@10 (%). Our model achieves the best results among models with fewer than 32B parameters. The left block reports evaluation with the robust verifier (syntax + functionality). The right block (w/o robust verifier) lacks functionality checks, leading to misleading correctness estimates. We denote our model trained from SFT with DeepSeek-R1 (resp. GPT-OSS 120B) as TritonRL (D) (resp. (G)) and report the results of SFT-only variants without RL for reference (See Section 2.1.1). We additionally denote our models trained with input augmentation (IA) as + IA, which uses RL datasets where diverse input shapes are augmented (See Section 2.1.2).
Model #Params Level1 (robust verifier) Level1 (w/o robust verifier)
valid compiled / correct $\textbf{fast}_{1}$ / $\textbf{fast}_{2}$ mean speedup compiled / correct
Qwen3 (base) 88B $99.0$ $99.0$ / $23.0$ $9.0$ / $5.0$ $0.41$ $100.0$ / $25.0$
Qwen3 1414B $100.0$ $99.0$ / $35.0$ $9.0$ / $7.0$ $0.48$ $100.0$ / $35.0$
Qwen3 3232B $100.0$ $100.0$ / $62.0$ $39.0$ / $16.0$ $0.93$ $100.0$ / $67.0$
KernelLLM 88B $42.0$ $40.0$ / $20.0$ $0.0$ / $0.0$ $0.05$ $99.0$ / $34.0$
AutoTriton 88B $97.0$ $92.0$ / $57.0$ $25.0$ / $10.0$ $0.95$ $100.0$ / $87.0$
TritonRL (D) 88B $100.0$ $\mathbf{100.0}$ / $78.0$ $36.0$ / $13.0$ $1.10$ $100.0$ / $81.0$
TritonRL (D) + IA 88B $100.0$ $99.0$ / $78.0$ $34.0$ / $15.0$ $1.05$ $100.0$ / $81.0$
   SFT w. DeepSeek-R1 (w.o. RL) 88B $100.0$ $\mathbf{100.0}$ / $54.0$ $28.0$ / $12.0$ $0.88$ $100.0$ / $60.0$
TritonRL (G) 88B $100.0$ $\mathbf{100.0}$ / $76.0$ $35.0$ / $17.0$ $1.09$ $100.0$ / $78.0$
TritonRL (G) + IA 88B $100.0$ $\mathbf{100.0}$ / $\mathbf{88.0}$ $\mathbf{41.0}$ / $\mathbf{22.0}$ $\mathbf{1.26}$ $100.0$ / $88.0$
   SFT w. GPT-OSS 120B (w.o. RL) 88B $100.0$ $\mathbf{100.0}$ / $64.0$ $26.0$ / $18.0$ $1.03$ $100.0$ / $73.0$
Claude-3.7 - $100.0$ $100.0$ / $63.0$ $29.0$ / $16.0$ $1.06$ $100.0$ / $73.0$
DeepSeek-R1 685685B $100.0$ $100.0$ / $69.0$ $28.0$ / $13.0$ $1.04$ $100.0$ / $74.0$
GPT-OSS 120120B $100.0$ $100.0$ / $81.0$ $33.0$ / $22.0$ $1.20$ $100.0$ / $87.0$
Model #Params Level2 (robust verifier) Level2 (w/o robust verifier)
valid compiled / correct $\textbf{fast}_{1}$ / $\textbf{fast}_{2}$ mean speedup compiled / correct
Qwen3 (base) 88B $38.0$ $37.0$ / $0.0$ $0.0$ / $0.0$ $0.00$ $100.0$ / $56.0$
Qwen3 1414B $40.0$ $38.0$ / $1.0$ $0.0$ / $0.0$ $0.00$ $100.0$ / $77.0$
Qwen3 3232B $36.0$ $35.0$ / $2.0$ $2.0$ / $1.0$ $0.10$ $100.0$ / $68.0$
KernelLLM 88B $3.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $0.00$ $100.0$ / $4.0$
AutoTriton 88B $11.0$ $10.0$ / $1.0$ $0.0$ / $0.0$ $0.00$ $100.0$ / $94.0$
TritonRL (D) 88B $63.0$ $63.0$ / $23.0$ $12.0$ / $3.0$ $0.32$ $100.0$ / $77.0$
TritonRL (D) + IA 88B $52.0$ $51.0$ / $24.0$ $16.0$ / $7.0$ $0.31$ $100.0$ / $79.0$
   SFT w. DeepSeek-R1 (w.o. RL) 88B $55.0$ $51.0$ / $13.0$ $10.0$ / $3.0$ $0.31$ $100.0$ / $67.0$
TritonRL (G) 88B $91.0$ $91.0$ / $23.0$ $15.0$ / $7.0$ $0.29$ $98.0$ / $30.0$
TritonRL (G) + IA 88B $\mathbf{93.0}$ $\mathbf{93.0}$ / $\mathbf{28.0}$ $\mathbf{19.0}$ / $\mathbf{10.0}$ $\mathbf{0.48}$ $96.0$ / $37.0$
   SFT w. GPT-OSS 120B (w.o. RL) 88B $88.0$ $88.0$ / $13.0$ $9.0$ / $3.0$ $0.21$ $97.0$ / $25.0$
Claude-3.7 - $34.0$ $34.0$ / $14.0$ $4.0$ / $1.0$ $0.10$ $100.0$ / $65.0$
DeepSeek-R1 685B $31.0$ $31.0$ / $14.0$ $9.0$ / $4.0$ $0.24$ $100.0$ / $79.0$
GPT-OSS 120B $39.0$ $39.0$ / $15.0$ $12.0$ / $4.0$ $0.25$ $100.0$ / $81.0$
@@ 3.1 Training and Evaluation Setups
We implement the training pipeline using the VeRL framework (Sheng et al., 2025) and initialize all experiments from Qwen3-8B (Team, 2025). Unless noted otherwise, RL training starts from the SFT checkpoint distilled from DeepSeek-R1 (denoted by TritonRL (D)) and uses the reward $r$ in (2) with $\alpha^{*}=0.1$. SFT is performed with batch size $16$ and learning rate $1\times 10^{-5}$; RL uses batch size $32$ and learning rate $1\times 10^{-6}$. We denote the model fine-tuned from the GPT-OSS 120120B SFT checkpoint as TritonRL (G); most training hyperparameters were kept the same (see Appendix E for details).
@@
We evaluate TritonRL on KernelBench (Li et al., 2025a) [^2] We use the Triton backend version of KernelBench from https://github.com/ScalingIntelligence/KernelBench/pull/35.. KernelBench offers an evaluation framework covering 250 tasks, divided into Level 1 (100 single-kernel tasks, such as convolution), Level 2 (100 simple fusion tasks, such as conv+bias+ReLU), and Level 3 (50 full architecture tasks, such as MobileNet), to assess LLM proficiency in generating efficient CUDA kernels. We conduct experiments mainly on the Level 1 and Level 2 tasks from KernelBench. The prompts used for these benchmarks are provided in Appendix I.1.
@@
We evaluate the performance of LLMs for generating Triton code in terms of (1) Validity (syntax and functionality); (2) Correctness (compilation and correct output); (3) Speedup (relative execution time improvement). We report fast1 and fast2 to indicate the model’s ability to generate Triton code that is at least as fast as or twice as fast as the reference PyTorch implementation, respectively. The formal definition of metrics is provided in Appendix C. We measure the pass@$k$ metrics for each aspect, which indicates the ratio of generating at least one successful solution among $k$ sampled attempts. We use $k=10$ as a default unless specified. We test both Triton codes and reference PyTorch codes on an NVIDIA L40S.
@@
We compare TritonRL with several baselines, including KernelLLM (Fisches et al., 2025) and AutoTriton (Li et al., 2025b), which are fine-tuned LLMs specifically for Triton programming. We also include our base model Qwen3-8B (Team, 2025) without any fine-tuning, fine-tuned Qwen3-8B only after SFT, and larger Qwen3 models (e.g., Qwen3-14B and Qwen3-32B). Additionally, we evaluate Claude-3.7 (Anthropic, 2025) with unknown model size and large model classes beyond 100B (e.g., GPT-OSS 120B (OpenAI, 2025), DeepSeek-R1-0528 (Guo et al., 2025)) as a reference.